# Ice Cream Doesn’t Cause Drowning: Benchmarking LLMs Against Statistical Pitfalls in Causal Inference

### v1

@@ Abstract
Reliable causal inference is essential for making decisions in high-stakes areas like medicine, economics, and public policy. However, it remains unclear whether large language models (LLMs) can handle rigorous and trustworthy statistical causal inference. Current benchmarks usually involve simplified tasks. For example, these tasks might only ask LLMs to identify semantic causal relationships or draw conclusions directly from raw data. As a result, models may overlook important statistical pitfalls, such as Simpson’s paradox or selection bias. This oversight limits the applicability of LLMs in the real world. To address these limitations, we propose CausalPitfalls, a comprehensive benchmark designed to rigorously evaluate the capability of LLMs in overcoming common causal inference pitfalls. Our benchmark features structured challenges across multiple difficulty levels, each paired with grading rubrics. This approach allows us to quantitatively measure both causal reasoning capabilities and the reliability of LLMs’ responses. We evaluate models using two protocols: (1) direct prompting, which assesses intrinsic causal reasoning, and (2) code-assisted prompting, where models generate executable code for explicit statistical analysis. Additionally, we validate the effectiveness of this judge by comparing its scoring with assessments from human experts. Our results reveal significant limitations in current LLMs when performing statistical causal inference. The CausalPitfalls benchmark provides essential guidance and quantitative metrics to advance the development of trustworthy causal reasoning systems.
@@ 1 Introduction
Causal inference [1, 2] is fundamental to decision-making across diverse fields. For instance, accurately determining the effectiveness and safety of a vaccine is pivotal in public health decisions [3]. However, identifying causal relationships with both reliability and interpretability remains challenging. In practice, individuals without formal statistical training frequently fall into subtle pitfalls, leading to plausible yet incorrect conclusions. A classic illustration is the erroneous conclusion that ice cream sales cause drowning incidents — overlooking the hidden confounder of hot weather causing both events [4, 5, 1].
Given these complexities, automated tools like large language models (LLMs) present promising avenues, demonstrated by their effectiveness in scientific problem-solving [6, 7] and clinical reasoning [8]. Recent studies [9, 10, 11] have evaluated LLMs’ abilities to evaluate accuracy in causal-effect estimation, but these benchmarks often neglect crucial aspects like robustness, interpretability, and susceptibility to common causal pitfalls. As a result, LLMs can produce seemingly convincing yet misleading causal conclusions.
To illustrate why reliability assessment is crucial, we highlight two significant pitfalls (detailed in Section 3). First, we find that LLMs frequently over-rely on prior knowledge rather than analyzing empirical data carefully. For example, given datasets on ice cream sales, temperature, and drowning incidents, whether causal or random, LLMs consistently identify temperature as a confounder without empirical justification. Second, LLMs can easily be misled by superficial semantic cues. For example, we evaluated whether LLMs could accurately determine if a drink positively impacts health. Despite using identical data, the LLM inferred a healthy effect when the drink was labeled “HealthPlus,” yet inferred an unhealthy effect when labeled “UltraSugar.” These examples demonstrate that without evaluation of reliability, models may produce convincing yet deeply flawed conclusions. Thus, accuracy alone provides an insufficient measure for assessing LLM-based causal inference, motivating the need for a targeted, comprehensive evaluation benchmark.
Figure 1: Overall Message: Our results reveal a clear reliability gap in causal inference when LLMs rely only on direct prompting, with all models struggling most on mediation and external validity questions. Introducing code assisted prompting leads to substantial gains across every task and brings all models closer together in performance. This shows that executable analysis is essential for large language models to handle complex statistical challenges and deliver trustworthy causal conclusions.
@@ 1.1 Main contributions
First, we introduce CausalPitfalls, a novel comprehensive benchmark specifically designed to evaluate the reliability of large language models (LLMs) in statistical causal inference. Unlike existing benchmarks primarily focused on accuracy, our benchmark explicitly targets model susceptibility to common causal pitfalls as shown in Figure 1, including (1) Confounding Biases and Spurious Associations, (2) Interventions and Experimental Reasoning, (3) Counterfactual Reasoning and Hypotheticals, (4) Mediation and Indirect Causal Effects, (5) Causal Discovery and Structure Learning, and (6) Causal Generalization and External Validity [12, 13]. These categories are structured into 15 distinct challenges, encompassing a total of 75 evaluation questions and 75 carefully constructed datasets that systematically test the robustness of LLM causal reasoning capabilities.
Second, we comprehensively evaluate the reliability of five leading LLMs under two distinct evaluation protocols: (1) direct prompting, assessing intrinsic causal reasoning from raw data, and (2) code-assisted prompting, where models generate executable code to perform explicit statistical analyses before responding. This dual-protocol approach provides a detailed quantitative assessment, highlighting areas where computational assistance significantly improves causal reasoning and where intuitive reasoning could suffice or even outperform explicit analyses.
@@ 1.2 Related Work
1. Causal Inference and Statistical Pitfalls. Causal inference from observational data is inherently challenging because counterfactuals are unobservable and confounding is ubiquitous [1, 2]. Causal inference methods for addressing confounders, whether the confounders are measured [14, 15, 16] or latent [17, 18], depend on restrictive model assumptions that are often difficult to verify empirically. Inferring causal direction similarly hinges on stringent structural equation assumptions [19, 20] or auxiliary information such as valid instruments [21]. Mediation analysis [22, 23], which targets specific causal pathways, demands careful adjustment for intermediate variables to avoid post-treatment bias. Finally, transporting causal conclusions across different domains requires justification of source-target invariances and methodologies for causal knowledge transfer [24, 25]. Rigorously confronting each of these challenges is essential to conduct a reliable causal analysis.
2. LLMs for Causal Reasoning. Recent studies have extensively investigated the causal reasoning capabilities of LLMs [26, 27, 28]. For example, Kiciman et al. [29] demonstrated that LLMs can infer causal relationships only from variable names, outperforming traditional statistical approaches [13]. However, these evaluations focus on scenarios involving commonsense causality. To bridge this gap, Jin et al. [30] introduced synthetic datasets generated from causal graphs, thereby enabling the assessment of LLMs’ causal reasoning performance in contexts extending beyond commonsense knowledge. Additionally, recent works [31, 32] have assessed LLMs’ capabilities in data-driven causal inference tasks, focusing on accuracy in estimating causal effects and recovering DAG structures from observational data.
@@ 2.1 Pitfall Categories and Challenges
Many of these scenarios can also be framed as purely statistical association problems. However, throughout this benchmark, we adopt a causal interpretation grounded in either the Neyman–Rubin potential outcomes framework [33, 34] or Pearl’s do-calculus [1]. This distinction ensures that our tasks explicitly target causal reasoning, rather than surface-level correlations.
One illustrative example of these pitfalls is Simpson’s paradox [35] (Fig. 2a), a commonly misunderstood statistical phenomenon. Simpson’s paradox occurs when overall data seem to suggest one conclusion: for example, patients recover better with a particular medical treatment. At first glance, the treatment might appear beneficial. However, once the patients are divided into subgroups based on age, the same data show that the treatment is harmful within each age group. An incorrect analysis would overlook the importance of these subgroups, mistakenly suggesting the treatment is effective. Importantly, the stratification variable in Simpson’s paradox is assumed to be a confounder, not a collider.
@@ 2.2 Evaluation Protocols and Data
Each challenge includes five questions with difficulty levels ranging from very easy to very hard (Table 2), allowing us to assess LLM performance across varying levels of complexity (Fig. 2b). Specifically, easier questions provide more guidance and suggestions, while harder questions offer fewer hints and thereby require LLMs to independently recognize and address potential pitfalls.
@@
Easy Evaluate whether {TREATMENT} causally affects {OUTCOME}. Consider the potential influence of the confounder ({CONFOUNDER}), adjust accordingly, and explicitly assess Simpson’s paradox. Provide adjusted effect sizes with uncertainty estimates and a recommendation.
@@
To construct datasets tailored to each challenge, we utilize causal graphs following Pearl et al. [12] and Peters et al. [13]. For every statistical pitfall, we select causal graphs that capture its unique complexities and characteristics. Each challenge is accompanied by five distinct datasets, each containing over 500 samples for comprehensive evaluation. Our simulation approach uses structural causal models based on directed acyclic graphs (DAGs), where each structural equation represents a causal mechanism rather than merely a statistical association. The coefficients in these equations directly encode the causal effects, allowing us to define the ground truth against which inference methods can be evaluated. This approach is mathematically equivalent to simulating potential outcomes under the specified causal structure.
@@ 2.3 Evaluation Metrics
To evaluate the reliability of LLMs for causal inference, we developed detailed grading rubrics for each causal pitfall, informed by guidelines from Sterne et al. [36], Vandenbrouckel et al. [37]. Each benchmark challenge includes multiple questions, each assigned points based on how effectively the model addresses the specific pitfall (see Appendix for detailed rubric). The total score for a challenge is the sum of points obtained across these questions, and max_score is the maximum achievable score. To enable fair comparisons across challenges, we compute a normalized score:
[…]
We evaluate LLM responses automatically using an independent GPT-4o model [7] to minimize potential biases. To validate the accuracy of this automated evaluation, we additionally engaged three statisticians to manually grade 150 randomly selected responses. We measure consistency between automated and human scores using the gap metric:
$\text{Gap}=\frac{1}{150}\sum_{i=1}^{150}\frac{|\text{score}_{\text{LLM}}^{(i)}-\text{score}_{\text{human}}^{(i)}|}{s_{\max,i}},$
where $s_{\max,i}$ is the maximum score of corresponding challenge, and $\text{score}_{\text{LLM}}^{(i)},~{}\text{score}_{\text{human}}^{(i)}\in\mathbb{N}^{+}$ are scores from automated and human evaluations, respectively.
@@
3 Motivating Failures: Why Accuracy Is Not Enough
When we assess causal inference using LLMs, high accuracy might initially seem sufficient. After all, if a model provides correct answers, why worry further? However, accuracy alone can create a dangerous illusion of reliability, masking vulnerabilities beneath confident and plausible conclusions.
To illustrate why our evaluation is necessary, we present two failure cases where LLMs make fundamentally incorrect causal inferences despite appearing to succeed. These examples, centered on confounding and semantic bias, highlight the limitations of current models and underscore the need for our benchmark’s focus on reliability, not just correctness. In real-world applications such as healthcare or policy, these pitfalls can lead to dangerously misleading decisions.
Table 3: LLMs rely purely on intuitive prior beliefs, ignoring actual data.
Dataset Type GPT-4o Claude-3.5-sonnet Gemini-2.0-flash
Real Confounder Data Correct (✓) Correct (✓) Correct (✓)
Random (Pure Noise) Data Incorrect ($\times$) Incorrect ($\times$) Incorrect ($\times$)
Over-reliance on Intuitive Prior Knowledge.
To understand this subtle yet impactful issue, imagine investigating a classical example often used to illustrate causal inference principles: the correlation between ice cream sales, temperature, and drowning incidents. Common intuition suggests a confounder: warmer temperatures encourage swimming, thereby increasing drowning incidents, and simultaneously encourage ice cream consumption. Inspired by this intuitive reasoning, we constructed two distinct datasets to evaluate LLM reasoning objectively:
  • Real Confounder Data: Temperature genuinely influences both ice cream sales and drowning incidents, forming a valid causal structure.
  • Random (Pure Noise) Data: Temperature, ice cream sales, and drowning incidents are independently generated, ensuring no actual causal relationships.
Surprisingly yet worryingly, LLMs consistently identified temperature as a confounder in both datasets. They confidently provided causal explanations based purely on intuitive prior beliefs, even when the data offered no empirical support (summarized in Table 3). Although in this simple scenario the intuitive conclusions aligned with common sense, this ignorance of empirical data raises severe risks. In real-world scenarios where intuitive reasoning might fail or mislead, this can dangerously reinforce incorrect causal interpretations. Thus, rigorous evaluation methods that test for data-driven reasoning are crucial.
@@ 3 Motivating Failures: Why Accuracy Is Not Enough
What happens if someone deliberately exploits this intuitive bias? To investigate this risk, we designed an adversarial scenario assessing whether a beverage has a beneficial or harmful impact on health. First, we generated synthetic datasets according to a realistic causal structure (Fig. 3). In these datasets, beverage consumption directly influenced health outcomes positively or negatively, while factors such as lifestyle and health awareness independently affected both beverage intake and health. Crucially, the beverage’s brand name itself had no genuine causal effect.
@@
However, by merely changing the beverage’s label from healthy-sounding brand “HealthPlus”’ to harmful-sounding brand “UltraSugar”, we observed a striking shift in LLM conclusions despite identical underlying data (Table 4). Specifically, GPT-4o and Gemini-2.0-flash consistently concluded the beverage labeled “HealthPlus” was beneficial and the one labeled “UltraSugar” harmful. This susceptibility highlights how easily LLM-based causal inference can be attacked by superficial semantic tricks. In real-world applications, where decisions often have substantial consequences, this vulnerability leads to serious risks. Thus, there is an urgent and practical need for reliable evaluation benchmarks designed to measure and enhance the resilience of causal reasoning against extensive causal inference pitfalls.
Table 4: Branding bias evaluation across LLMs. A checkmark (✓) indicates a beneficial effect; a cross ($\times$) indicates a harmful effect. Model conclusions matching the true effect indicate correct inference.
True Effect Brand Name GPT-4o Gemini-2.0-flash Claude-3.5-sonnet
✓(Beneficial) HealthPlus (✓)
UltraSugar ($\times$) $\times$ $\times$
$\times$ (Harmful) HealthPlus (✓)
UltraSugar ($\times$) $\times$ $\times$

### v2

@@ Abstract
Reliable causal inference is essential for making decisions in high-stakes areas like medicine, economics, and public policy. However, it remains unclear whether large language models (LLMs) can handle rigorous and trustworthy statistical causal inference. Current benchmarks usually involve simplified tasks. For example, these tasks might only ask LLMs to identify semantic causal relationships or draw conclusions directly from raw data. As a result, models may overlook important statistical pitfalls, such as Simpson’s paradox or selection bias. This oversight limits the applicability of LLMs in the real world. To address these limitations, we propose CausalPitfalls, a comprehensive benchmark designed to rigorously evaluate the capability of LLMs in overcoming common causal inference pitfalls. Our benchmark features structured challenges across multiple difficulty levels, each paired with grading rubrics. This approach allows us to quantitatively measure both causal reasoning capabilities and the reliability of LLMs’ responses. We evaluate models using two protocols: (1) direct prompting, which assesses intrinsic causal reasoning, and (2) code-assisted prompting, where models generate executable code for statistical analysis. Additionally, we validate the effectiveness of this judge by comparing its scoring with assessments from human experts. Our results reveal significant limitations in current LLMs when performing statistical causal inference. The CausalPitfalls benchmark provides essential guidance and quantitative metrics to advance the development of trustworthy causal reasoning systems. Our code is publicly available at CausalPitfalls.
@@ 1 Introduction
Causal inference (Pearl, 2009; Imbens and Rubin, 2015) is fundamental to decision-making across diverse fields. For instance, accurately determining the effectiveness and safety of a vaccine is pivotal in public health decisions (Voysey et al., 2021). However, identifying causal relationships with both reliability and interpretability remains challenging. In practice, individuals without formal statistical training frequently fall into subtle pitfalls, leading to plausible yet incorrect conclusions. A classic illustration is the erroneous conclusion that ice cream sales cause drowning incidents — overlooking the hidden confounder of hot weather causing both events (Pearl, 2009; Greenland and Robins, 1986; Rosenbaum, 1987).
Given these complexities, automated tools like large language models (LLMs) present promising avenues, demonstrated by their effectiveness in scientific problem-solving (Lewkowycz et al., 2022; Achiam et al., 2023) and clinical reasoning (Singhal et al., 2023). Recent studies (Wang, 2024; Dhawan et al., 2024; Liu et al., 2024) have evaluated LLMs’ abilities to evaluate accuracy in causal-effect estimation, but these benchmarks often neglect crucial aspects like robustness, interpretability, and susceptibility to common causal pitfalls. As a result, LLMs can produce seemingly convincing yet misleading causal conclusions.
To illustrate why reliability assessment is crucial, we highlight two representative failure modes (detailed in Section 3). First, LLMs can ignore strong data evidence, but in favor of superficial semantic cues: in a synthetic health scenario with identical datasets, LLMs concluded the drink was beneficial when labeled “HealthPlus” and harmful when labeled “UltraSugar,” even when the data indicated the opposite. Second, LLMs can mistake random variation for genuine causal structure: when tested on research funding data from the Netherlands, LLMs attributed differences in success rates to gender bias or Simpson’s paradox, despite statistical analyses showing that neither claim is supported. These cases demonstrate that LLMs may produce confident causal claims directly contradicted by the data, showing the need for benchmarks that assess causal inference reliability.
Figure 1: Overall Message: Our results reveal a clear reliability gap in causal inference when LLMs rely only on direct prompting, with all models struggling most on mediation and external validity questions. Introducing code-assisted prompting leads to substantial gains across every task and brings all models closer together in performance. This shows that executable analysis is essential for large language models to handle complex statistical challenges and deliver trustworthy causal conclusions. Full results for all evaluated LLMs are provided in Table 4.
@@ 1.1 Main contributions
First, we introduce CausalPitfalls, a novel comprehensive benchmark specifically designed to evaluate the reliability of large language models (LLMs) in statistical causal inference. Unlike existing benchmarks primarily focused on accuracy, our benchmark targets model susceptibility to common causal pitfalls as shown in Figure 1, including (1) Confounding Biases and Spurious Associations, (2) Interventions and Experimental Reasoning, (3) Counterfactual Reasoning and Hypotheticals, (4) Mediation and Indirect Causal Effects, (5) Causal Discovery and Structure Learning, and (6) Causal Generalization and External Validity (Pearl and others, 2003; Peters et al., 2017). These categories are structured into 15 distinct challenges, encompassing a total of 75 evaluation questions and 75 carefully constructed datasets that systematically test the robustness of LLM causal reasoning capabilities.
Second, we comprehensively evaluate the reliability of ten LLMs under two distinct evaluation protocols: (1) direct prompting, assessing intrinsic causal reasoning from raw data, and (2) code-assisted prompting, where models generate executable code to perform statistical analyses before responding. This dual-protocol approach provides a detailed quantitative assessment, highlighting areas where computational assistance significantly improves causal reasoning and where intuitive reasoning could suffice.
@@ 1.2 Related Work
1. Causal Inference and Statistical Pitfalls. Causal inference from observational data is inherently challenging because counterfactuals are unobservable and confounding is ubiquitous (Pearl, 2009; Imbens and Rubin, 2015). Causal inference methods for addressing confounders, whether the confounders are measured (Chan et al., 2016; Lin et al., 2023; Doss et al., 2024) or latent (Kang et al., 2016; Guo et al., 2022), depend on restrictive model assumptions that are often difficult to verify empirically. Inferring causal direction similarly hinges on stringent structural equation assumptions (Peters and Bühlmann, 2014; Li et al., 2024) or auxiliary information such as valid instruments (Chen et al., 2024). Mediation analysis (MacKinnon, 2012; Yang et al., 2021), which targets specific causal pathways, demands careful adjustment for intermediate variables to avoid post‐treatment bias. Finally, transporting causal conclusions across different domains requires justification of source-target invariances and methodologies for causal knowledge transfer (Wei et al., 2023; Chen et al., 2025). Rigorously confronting each of these challenges is essential to conduct a reliable causal analysis.
2. LLMs for Causal Reasoning. Recent studies have extensively investigated the causal reasoning capabilities of LLMs (Willig et al., 2022; Zečević et al., 2023; Qin et al., 2019). For example, Kiciman et al. (2023) demonstrated that LLMs can infer causal relationships only from variable names, outperforming traditional statistical approaches (Peters et al., 2017). However, these evaluations focus on scenarios involving commonsense causality. To bridge this gap, Jin et al. (2023) introduced synthetic datasets generated from causal graphs, thereby enabling the assessment of LLMs’ causal reasoning performance in contexts extending beyond commonsense knowledge. Additionally, recent works (Long et al., 2023; Zhou et al., 2024) have assessed LLMs’ capabilities in data-driven causal inference tasks, focusing on accuracy in estimating causal effects and recovering DAG structures from observational data.
@@ 2.1 Pitfall Categories and Challenges
Many of these scenarios can also be framed as purely statistical association problems. However, throughout this benchmark, we adopt a causal interpretation grounded in either the Neyman–Rubin potential outcomes framework (Splawa-Neyman et al., 1990; Rubin, 1974) or Pearl’s do-calculus (Pearl, 2009). This distinction ensures that our tasks target causal reasoning, rather than surface-level correlations.
One illustrative example of these pitfalls is Simpson’s paradox (Simpson, 1951) (Fig. 2a), a commonly misunderstood statistical phenomenon. Simpson’s paradox occurs when overall data seem to suggest one conclusion: for example, patients recover better with a particular medical treatment. At first glance, the treatment might appear beneficial. However, once the patients are divided into subgroups based on age, the same data show that the treatment is harmful within each age group. An incorrect analysis would overlook the importance of these subgroups, mistakenly suggesting the treatment is effective. Importantly, the stratification variable in Simpson’s paradox is assumed to be a confounder, not a collider.
@@ 2.2 Evaluation Protocols and Data
Each challenge includes five versions of the core question, ranging from very easy to very hard (Table 2). The easier versions give the model more guidance. For example, pointing out the confounder to adjust for or asking directly about Simpson’s paradox. As the difficulty increases, these hints are gradually removed. This setup lets us test whether models can still recognize and handle causal pitfalls when less direction is given (Fig. 2b).
@@
Easy Evaluate whether {TREATMENT} causally affects {OUTCOME}. Consider the potential influence of the confounder ({CONFOUNDER}), adjust accordingly, and assess Simpson’s paradox. Provide adjusted effect sizes with uncertainty estimates and a recommendation.
@@
To construct datasets tailored to each challenge, we utilize causal graphs following Pearl and others (2003) and Peters et al. (2017). For every statistical pitfall, we select causal graphs that capture its unique complexities and characteristics. Each challenge is accompanied by five distinct datasets, each containing over 500 samples for comprehensive evaluation. Our simulation approach uses structural causal models based on directed acyclic graphs (DAGs), where each structural equation represents a causal mechanism rather than merely a statistical association. The coefficients in these equations directly encode the causal effects, allowing us to define the ground truth against which inference methods can be evaluated. This approach is mathematically equivalent to simulating potential outcomes under the specified causal structure. The structural equations include both linear and non-linear forms (e.g., non-linear link functions and interaction terms), so the evaluation is not limited to purely linear relationships.
@@ 2.3 Evaluation Metrics
To evaluate the reliability of LLMs for causal inference, we developed detailed grading rubrics for each causal pitfall, informed by guidelines from Sterne et al. (2016); Vandenbrouckel et al. (2007). Each benchmark challenge includes multiple questions, each assigned points based on how effectively the model addresses the specific pitfall (see Appendix for detailed rubric). The total score for a challenge is the sum of points obtained across these questions, and max_score is the maximum achievable score. To enable fair comparisons across challenges, we compute a normalized score:
$\text{Normalized Score (\%)}=\frac{\text{score}}{\text{max\_score}}\times 100\%.$ (1)
We evaluate LLM responses automatically using an independent GPT-4o model (Achiam et al., 2023) to minimize potential biases. To validate the accuracy of this automated evaluation, we additionally engaged three statisticians to manually grade 150 randomly selected responses. We measure consistency between automated and human scores using the gap metric:
$\text{Gap}=\frac{1}{150}\sum_{i=1}^{150}\frac{|\text{score}_{\text{LLM}}^{(i)}-\text{score}_{\text{human}}^{(i)}|}{s_{\max,i}}\in[0,1],$
where $s_{\max,i}$ is the maximum score of corresponding challenge, and $\text{score}_{\text{LLM}}^{(i)},~\text{score}_{\text{human}}^{(i)}\in\mathbb{N}^{+}$ are scores from automated and human evaluations, respectively. Here, the gap metric ranges from 0 to 1, where 0 indicates perfect agreement.
@@
3 Illustrative Pitfalls in Causal Reasoning
When evaluating causal inference with LLMs, surface-level answers can create an illusion of competence. Statistical causal inference requires grounding conclusions in evidence, checking assumptions, and ruling out alternatives. LLMs, however, may produce confident but flawed outputs that rely on irrelevant cues or statistical artifacts, giving a false sense of robustness.
We illustrate this problem with two failure cases. The first shows how models can base conclusions on superficial semantic cues instead of data, while the second shows how they may misinterpret random variation as causal structure.
@@ 3 Illustrative Pitfalls in Causal Reasoning
As an illustrative example, we examined whether LLMs can be misled by superficial cues when drawing causal conclusions. We constructed a synthetic scenario in which beverage consumption affected health outcomes. In the setting, other factors, lifestyle and health awareness, acted as confounders, but the brand label itself had no causal role (Fig. 3).
@@
Despite identical data, simply changing the brand name from a healthy-sounding label (HealthPlus) to a harmful-sounding one (UltraSugar) induced changes in LLMs’ conclusions (Table 3). GPT-4o and Gemini-2.0-flash, for example, made a conclusion aligning purely with the brand semantics, rather than the provided data.
The branding bias example shows that LLMs may rely on superficial semantic cues, attributing causal effects to labels even when the underlying data provide evidence to the contrary.
Table 3: Branding bias evaluation. Each row represents a combination of the beverage’s label (“HealthPlus” or “UltraSugar”) and the true effect of the given data (beneficial or harmful). A checkmark (✓) indicates a beneficial conclusion and a cross ($\times$) a harmful conclusion; correct inferences are those that match the true effect.
Brand Label True Effect (Data) GPT-4o Gemini-2.0-flash Claude-3.5-sonnet
HealthPlus
UltraSugar $\times$ $\times$
HealthPlus $\times$
UltraSugar $\times$ $\times$ $\times$
Spurious Causal Inference from Random Patterns.
As a second illustrative example, we show how LLMs can mistake random variation in real-world data for genuine causal structure. Specifically, we evaluated LLMs on real data from a PNAS study (Van der Lee and Ellemers, 2015) that analyzed funding success rates across academic disciplines in the Netherlands. When asked if the data reveal gender bias favoring male applicants, all of the tested LLMs drew incorrect conclusions, either attributing the differences in raw percentages directly to gender or considering Simpson’s paradox.
However, a careful statistical analysis as illustrated in the Appendix following Irizarry (2019) shows that women are not disproportionately applying to more competitive disciplines, so Simpson’s paradox does not apply. Furthermore, computing log-odds ratios divided by their standard errors across disciplines and adjusting for multiple comparisons reveals no statistical evidence of gender bias within any department. None of the tested LLMs recognized these key insights.
The spurious inference example shows that LLMs may mistake random variation in observational data for genuine causal structure, failing to apply the statistical checks needed to rule out noise.